The x-coordinate of a particle moving along x-axis varies with time as x = 6 - 4t + t^2, where x is in meter and t is in second. Then distance covered by the particle in first 3 seconds is :
  • 1 m
  • 3 m 
  • 5 m 
  • 7 m
 
I got the answer as 3 m but the correct answer is given as 5 m . I am confused , so please tell the correct answer with explanation !

Grade:12

3 Answers

Adarsh
6 years ago
Dear Harish ..
 
At t=0 particle is at (6,0).
At t=1 , particle is at  ( 3,0) 
At t=2 ,  particle is at (2,0)
At t=3 , particle is at (3,0) 
So the distance travelled is (6 – 3) + (3 – 2) + (3 – 2) = 5m..
Thanks and Regards 
Adarsh
Harish
6 years ago
Sorry , but I didn’t understood why you didi this :
”So the distance travelled is (6 – 3) + (3 – 2) + (3 – 2) = 5m..”
 
Adarsh
6 years ago
During 1st second its coordinate changes from (6,0) to (3,0)  so distance travelled is 3m ..
in 2nd second its coordinate changes from (3,0) to (2,0) and distance travelled is 1m ..
in 3rd second its coordinate changes from (2,0) to (3,0) and distance travelles is 1m ..
so total distance in 3 second = 3 + 1 + 1 = 5m .. 
Hope this will help you..